# DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought

### v2

@@ 3.1 Experimental Setups
And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery. {CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
@@
This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought. {CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
@@
I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea. {CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
@@
The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky. {CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
@@
3.2 Comparison Models
We use vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as the comparison models. Besides, two O1-like LLMs, i.e., Marco-o1-7B and QwQ-32B-preview, are also conducted as baselines.
3.3 Main Results
Table 2 shows the experimental results. After instruction tuning on the long-thought MT training data, DRT-series LLMs significantly outperform the corresponding vanilla backbones. For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 4.04 GRF, 1.42 CometKiwi, 4.46 GRB, 8.52 BLEU and 3.41 CometScore. DRT-o1-8B outperforms Llama-3.1-8B-Instruct by 5.24 GRF, 0.71 CometKiwi, 7.50 GRB, 14.12 BLEU and 4.23 CometScore. Besides, DRT-o1-14B outperforms all baselines (including QwQ-32B-preview and Marco-o1-7B) in terms of all metrics, showing the effectiveness of long thought in MT.
@@ 3.4 Case Study
Table 3 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview and DRT-o1-14B. With the help of long thought, the translations of DRT-o1-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality. Besides, in addition to DRT-o1-14B, some translation snippets of other LLMs can also show a great performance (marked in green). This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate the capability.
Figure 4: Time cost during inference on the testing set.
3.5 Inference Time Analysis
When evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation. Specifically, we use a single NVIDIA A100 GPU (40G) to deploy and call each LLM, and calculate the corresponding generation time in the test set. As shown in Figure 4, the average time costs of DRT-o1 models are significantly higher than vanilla LLMs. This is because DRT-o1 models should first generate the long thought and then provide the final translation, thus needing more inference time. This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
@@ 4 Related Work
4.1 O1-like LLMs
Recently, O1-like LLMs have shown great performance in reasoning tasks, especially math and coding tasks. After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1. For example, Qin et al. (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process. With only 327 training samples, the journey learning enhances LLMs’ long-thought ability. Huang et al. (2024) explore the data distillation from existing O1-like models, and show the effectiveness of data distillation. Zhang et al. (2024) leverage Monte Carlo Tree Search (MCTS) to synthesize reasoning-enhanced code data, and train O1-Coder. Marco-o1 Zhao et al. (2024) is proposed to deal with open-ended text generation. Marco-o1 shows the effectiveness of a MCTS-enhanced inference method, and achieves great performance on math problems. Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
4.2 Literature Translation
Different from translating standard MT corpora (e.g., news articles), translating literature books is more difficult since it often requires equivalence beyond the word level Thai et al. (2022). Besides, it is also difficult to evaluate literature translation using automatic metrics, and previous literature translation work typically relies on human evaluation Fonteyne et al. (2020); Karpinska and Iyyer (2023). Due to its difficulty, early work is limited to small-scale attempts Genzel et al. (2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al. (2018). Recently, Karpinska and Iyyer (2023) utilize LLMs to perform literature translation, and show that discourse-level LLM translators achieve better performances compared with sentence-level approaches. Thai et al. (2022) introduce Par3 to benchmark LLMs’ literature translation capability from non-English languages to English. Wang et al. (2024b) propose DelTA, an online document-level MT agent to translate book-level long documents based on short-term memory, long-term memory, summaries and proper noun records.
@@ 5 Conclusion
In this paper, we introduce DRT-o1, an attempt to bring the success of long-thought reasoning to neural machine translation (MT). Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation. To collect the source sentences that are suitable for translation via long thought, we mine sentences with similes or metaphors from existing literature books. To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively. GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework. Based on the synthesized data, we train DRT-o1 models (using Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.
Appendix A GPT-4o Evaluator
@@
For GRB and GRF, we prompt GPT-4o (2024-08-06 version) as the MT evaluator in the reference-based and reference-free manners, respectively. The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation (c.f. Figure 5).
Appendix B Complete Generation of DRT-o1
Figure 6 shows an example of DRT-o1-14B. As we can see, the model learns the thought process in our data collection. DRT-o1-14B first performs the word-level translation, and then attempts the preliminary translation. Next, it iteratively refines its translation until it thinks the translation is good enough.

### v3

@@ 3.5 Inference Time Analysis
And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery. {CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
@@
This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought. {CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
@@
I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea. {CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
@@
The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky. {CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
@@ 3.6 Case Study
Table 5 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview, Qwen2.5-14B-SFT and DRT-14B. With the help of long thought, the translations of DRT-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality. In addition to DRT-14B, some translation snippets of other LLMs can also show a great performance (marked in green). This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate this capability.
@@ 4 Related Work
O1-like LLMs. Recently, O1-like LLMs have shown great performance in reasoning tasks, especially math and coding tasks. After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1. For example, Qin et al. (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process. Huang et al. (2024) explore the data distillation from existing O1-like models, and show the effectiveness of data distillation. Zhang et al. (2024) leverage Monte Carlo Tree Search (MCTS) to synthesize reasoning-enhanced code data, and train O1-Coder. Marco-o1 Zhao et al. (2024) is proposed to deal with open-ended text generation. More recently, DeepSeek-R1 Guo et al. (2025) and Kimi K1.5 Team et al. (2025) are proposed, and show their promising reasoning ability.
Literature Translation. Different from translating standard MT corpora (e.g., news articles), translating literature books is more difficult since it often requires equivalence beyond the word level Thai et al. (2022). Besides, it is also difficult to evaluate literature translation using automatic metrics, and previous literature translation work typically relies on human evaluation Fonteyne et al. (2020); Karpinska and Iyyer (2023). Due to its difficulty, early work is limited to small-scale attempts Genzel et al. (2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al. (2018). Recently, Karpinska and Iyyer (2023) utilize LLMs to perform literature translation, and show that discourse-level LLM translators achieve better performances compared with sentence-level approaches. Thai et al. (2022) introduce Par3 to benchmark LLMs’ literature translation capability from non-English languages to English.
@@ 5 Conclusion
In this paper, we introduce DRT, an attempt to bring the success of long-thought reasoning to neural machine translation (MT). Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation. To collect the source sentences that are suitable for translation via long thought, we mine sentences with similes or metaphors from existing literature books. To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively. Based on the synthesized data, we train DRT models. Extensive experiments on literature translation demonstrate the effectiveness of DRT models in terms of automatic evaluation. Case study and human evaluation further verify their superiority.
Limitations
While we show the effectiveness of long thought in MT, there are some limitations worth noting: (1) We focus on English-to-Chinese translation in this work, and future work could extend the data and the method to other translation directions. (2) There is still a lack of accurate automatic evaluation metrics for literary translation. Previous literature translation work typically relies on human evaluation Fonteyne et al. (2020); Karpinska and Iyyer (2023), and points out that BLEU and Comet might not be suitable for evaluating literature translation Karpinska and Iyyer (2023). This is because literary translations carry the responsibility of both semantic and critical interpretation, as they must address the challenge of achieving equivalence that often extends beyond the level of individual words Thai et al. (2022).
Ethical Considerations
We discuss the main ethical considerations of DRT models as follows: (1) Copyright. We mine literature sentences from 400 English books provided by the Project Gutenberg public-domain book repository [^5] https://www.gutenberg.org/, where the books are typically more than fifty years old and their copyrights have expired. The book data also has been extracted and released by Kryscinski et al. (2022). Therefore, we can construct DRT data based on these books, and further release our synthesized data. (2) Licenses. We will release our model checkpoints and synthesized data under CC-BY-NC-SA 4.0 license.
Appendix A Prompt in Data Synthesis
A.1 Prompts in Literature Book Mining
SYSTEM PROMPT: You are assigned to translate an English literary work into Chinese. The text may include descriptions or expressions that embody English cultural nuances, which may not resonate with Chinese language habits. In such instances, a literal translation may not be appropriate; instead, these sentences should be paraphrased to convey their intended meaning effectively. USER PROMPT: The English sentence is provided as follows: <english sentence> {sentence} </english sentence> Please begin by assessing whether the English sentence contains any metaphors or similes. If there are none, respond with "no metaphors and no similes." If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand. If it is suitable, format your response as follows (two lines): "your literal translation for metaphors/similes here (in Chinese)" "suitable" If it is unsuitable, please provide the reason for the unsuitability. Format your response as follows (three lines): "your literal translation for metaphors/similes here (in Chinese)" "unsuitable" "reason for unsuitability here (in Chinese)"
A.2 Prompts in Multi-Agent Framework
Translator Agent (Word-level translation)
Given an English sentence, identify the important words (usually nouns, verbs, technical terms, and named entities that require special attention in translation) and translate them into Chinese. Output the translations in JSON format, for example: {"EnglishWord1": "ChineseTranslation", "EnglishWord2": "ChineseTranslation"} The Chinese translations can be a single translation or multiple options as deemed appropriate.
Translator Agent (Preliminary translation)
SYSTEM PROMPT: Given an English sentence and a JSON object containing potential translations of important keywords, produce a Chinese literal translation of the entire sentence. Please directly output the Chinese translation without any descriptions. USER PROMPT: <English Sentence> {sentence} </English Sentence> <Potential Keyword Translation> {keyword translation} </Potential Keyword Translation>
Translator Agent (Refinement translation)
In the refine loop, the translator agent receives the feedback of the previous translation, and then provides a new translation. The prompt is a multi-turn dialogue between the translator and advisor, where the system prompt is the same as the preliminary translation.
Advisor Agent
Please rate the Chinese translation of the following English text and provide your comments and suggestions.
Evaluator Agent
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int}
A.3 Prompts in Thought Reformulation
A student is engaged in the task of translating an English sentence into Chinese. The English sentence is as follows: <English Sentence> {sentence} </English Sentence> This student constantly thinks about and optimizes his translation. The whole process is shown as follows: <Translation Process> {translation process} </Translation Process> Please polish the whole translation process into a long first-person self-reflection description (use the present tense). The self-reflection should begin with selecting the keywords from the English sentence, translating the keywords, and then attempt to translate the whole sentence, and then think about whether the translation is good or not, and iteratively make translation attempts. Finally, make a final translation decision. Output the self-reflection description directly without any additional descriptions or explanations. Each line in the self-reflection description can be regarded as a reasoning step to the translation.
Appendix B Details of Human Annotation
In Section 2.4, we employ human annotation to provide the quality comparison between two translations for a source sentence. Specifically, we employ three Chinese master students with high levels of fluency in both English and Chinese as our human annotators. For each sample, we give the source sentence and its two translation (without the scores provided by our evaluator agent) to all three annotators, and every annotator should provide one of the following judgments: (1) the first translation is better than the second one; (2) the second translation is better than the first one; (3) two translation are similar in quality. During annotation, we encourage the annotators to give differentiated judgments, i.e., judgment (1) or judgment (2). If three annotators give the same judgment for a sample, the judgment will be regarded as the final label. Otherwise, its label will be decided by a group meeting of all three annotators and a senior data scientist.
Appendix C GPT-4o Evaluator
@@
For GRB and GRF, we prompt GPT-4o (2024-08-06 version) as the MT evaluator in the reference-based and reference-free manners, respectively. The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation.
GRB Prompt:
Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese human reference: {ref} Chinese translation: {hyp} Score:
GRF Prompt:
Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese translation: {hyp} Score:
Model reference-free reference-based
GEA GRF CometKiwi GRB BLEU CometScore
Commercial LLMs
GPT-4o 71.88 85.57 73.01 82.78 34.51 79.41
o1-preview 78.01 87.11 73.70 83.86 30.65 80.12
DRT
DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
Table 6: Experimental results of comparing DRT with commercial LLMs. The bold and the underline denote the best and second-best performances, respectively.
Appendix D Implementation Details.
Automatic Evaluation. To calculate CometKiwi and CometScore, we leverage the official codes [^6] https://github.com/Unbabel/COMET and the official models [^7] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da. To calculate the BLEU score, we use the sacrebleu toolkit [^8] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
Training Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. All LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size. We use the DeepSpeed ZeRO-3 optimization Rasley et al. (2020). Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
Inference Details. When evaluating model performance on the test set, we use vLLM toolkit Kwon et al. (2023) to accelerate the model generation. We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05. For DeepSeek-R1 series (DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B), we follow the instruction [^9] https://github.com/deepseek-ai/DeepSeek-R1 to enforce them to avoid blank thinking. All experimental results listed in this paper are the average of 3 runs.
Appendix E Comparison with Commercial LLMs
As shown in Table 6, DRT-14B achieves competitive results with o1-preview, showing its superiority. Additionally, we observe that o1-preview significantly outperforms GPT-4o in terms of GEA. This finding highlights the effectiveness of long thought in machine translation. When applied to appropriate translation contexts, long thought can further enhance the authenticity of translations.